# Detection and Characterization of Coordinated Online Behavior: A Survey

### v1

@@ 2.6.1. Taxonomy
Figure 3 demonstrates the generality of the conceptual framework by presenting a taxonomy of coordinated online behavior based on the dimensions of authenticity and harmfulness. This analysis takes inspiration from the well-known conceptual framework of information disorder that categorizes instances of mis-, mal-, and disinformation based on the falseness and harmfulness of the information [136]. As shown in Figure 3, the dimensions of authenticity and harmfulness allow distinguishing between different types of coordinated online behavior, ranging from potentially nefarious to neutral and beneficial ones. Among the most problematic instances of online coordination are harmful and inauthentic phenomena such as information operations and disinformation campaigns, which are often initiated by malicious social bots and state-sponsored trolls [131, 19, 2]. Other problematic phenomena are those featuring harmful yet authentic behaviors, such as the activity of conspiracy theorists and hateful extremists [94, 91, 84]. Conversely, the activity of grassroots social movements, fandoms, and other activists represent instances of harmless and authentic coordination [119]. Finally, harmless yet inauthentic behaviors lay outside of the partially overlapping sets, and are exemplified by hacktivists, dissidents, and other anonymous protesters [55]. Alternative taxonomies can be obtained by leveraging the dimensions of orchestration and time-variance, which would highlight additional phenomena to those shown in Figure 3.
@@ 3. Problem Definition
The problem of identifying and investigating different types of coordinated online behavior involves defining two functions $f(\cdot)$ and $g(\cdot)$ that respectively implement the tasks of coordinated behavior detection and characterization, as outlined in Figure 4. Given a set of users and their actions on one or more online platforms, $f(\cdot)$ identifies possible coordinated groups of users. Instead, $g(\cdot)$ extracts additional information for each detected group, thus contributing to determining the nature, intent, and the overall characteristics of the involved actors (e.g., whether they are inauthentic, harmful, etc.). The detection and characterization tasks are related to the Definition 2.4 of coordinated online behavior and its components in that the function $f(\cdot)$ implementing the detection task does so via the analysis of user actions, while the function $g(\cdot)$ implementing the characterization task provides information about the actors and their intent. A comprehensive overview of the entire process is depicted in Figure 4. The input is represented by the set of users $U$ to analyze and their activities $H$. The detection task differentiates coordinated users from non-coordinated ones. Depending on the detection method, the distinction between the two can be expressed as binary labels assigned to the users, as two or more sets (e.g., clusters) of either coordinated or non-coordinated users, or as two or more coordinated or non-coordinated communities (i.e., nodes and edges) from a network. These are subsequently scrutinized during the characterization task, which computes a set of indicators for each coordinated user, set, or community. The indicators are selected so as to provide information about the characteristics of the coordinated actors and their behavior. For example, computing bot scores is a common method to estimate the inauthenticity of coordinated users.
@@ 3.1.2. Methods and output
(1) $f(U,H)=\begin{cases}\begin{aligned} &P=\{P_{1},\ldots,P_{i},\ldots P_{k}\},&&P_{i}=(V_{i},E_{i}),V_{i}\subseteq U&&\text{communities}\\ &C=\{C_{1},\ldots,C_{i},\ldots C_{k}\},&&C_{i}\subseteq U&&\text{clusters}\\ &B=\{B_{c},B_{u}\},&&B_{c}\cup B_{u}=U&&\text{binary labels}\\ \end{aligned}\end{cases}$
@@ 4.1.1. User selection
Multiple choices can be made to select a subset of relevant users. A common choice is to select the most active users as they produce the largest share of actions and content. Most active users can be defined as those who publish a large number of original posts (super-producers[67, 74, 102], or as those having a large number of re-shares (super-spreaders[99, 20, 56, 121, 70, 29, 88]. Other than activity, network centrality or influence, suspicious behavior, location, and timings are used for user selection [11, 73]. Furthermore, many other general criteria can also be adopted, such as selecting all users who posted certain keywords, or all followers of a given user. Combining multiple criteria allows for even more fine-grained user selections.
@@
This step builds a coordination network between the users $U^{\prime}\subseteq U$ previously selected. [^3] A few works follow the general approach of network science methods, but build networks of content rather than users. These are discussed in Section 4.1.5. A coordination network is a type of network where links exist only between coordinated nodes. As per Definition 2.4, coordination implies synergic actions between users. In network science methods, this concept is operationalized with co-actions. A co-action represents two users performing the same action on the same target or content. For instance, two users who comment, like, or re-share the same post are generating a co-action. As shown in Figure 6, the two coordinated users need not be directly connected in the social or interaction network. This characteristic makes coordination networks particularly suitable for surfacing coordination between seemingly unrelated users, such as those involved in inauthentic or harmful behavior, so much so that some scholars specifically refer to latent coordination networks [139]. In the most general case, a coordination network is a multiplex network $G(V,E,W,\mathcal{L})$. In order to build $G$, one must define the types of co-actions $C_{a}$ to consider (e.g., re-shares, mentions, follows, etc.). When multiple co-actions are used, $G$ is a multiplex network with $L$ layers, where =$\mathcal{L}=\{1,\dots,L\}$ is the set of layers, each corresponding to a co-action $i\in C_{a}$. However, the majority of existing works leverage a single co-action. In this case, the number of layers is $L=1$ and $G$ is a single layer network. Each layer in $G$ is an undirected weighted graph $G^{i}(V^{i},E^{i},W^{i})$, where $V^{i}\subseteq U^{\prime}$, $E^{i}$ and $W^{i}$ respectively denote the set of nodes, edges, and weights of layer $i$. We highlight that $V=\bigcup_{i\in\mathcal{L}}V^{i}$, $E=\bigcup_{i\in\mathcal{L}}E^{i}$, and $W=\bigcup_{i\in\mathcal{L}}W^{i}$. Given a layer $i$, an edge $e^{i}_{jk}$ is created if there exists a co-action of type $i$ between two users $(u_{j},u_{k})$. The edge weight $w^{i}_{jk}=sim_{i}(u_{j},u_{k})$ is obtained via a similarity function that computes pairwise user similarities in $U^{\prime}$. Different similarity functions can be used for different co-actions.
[…]
@@
reference action similarity filters community detection [19] retweet cardinality threshold, ADJ modularity clustering [51] retweet cardinality EDO Louvain [66] retweet cardinality threshold, ADO Louvain [68] retweet cardinality backbone, ADJ Louvain [109] retweet cardinality EDO [99, 20, 56, 70, 29, 121] retweet cosine similarity TF-IDF backbone Louvain [122] retweet cosine similarity TF-IDF backbone, EDO Leiden [137] retweet cardinality threshold, ADJ [18] retweet, tweet cardinality threshold, EDO Louvain, connected components [63] retweet, tweet cardinality threshold, EDO Louvain [110] retweet, tweet cardinality threshold, EDO [27] tweet cardinality ADO [67] tweet cardinality threshold cohesive campaign [101] tweet text similarity threshold, ADO [95, 93] parley cardinality threshold, kNN graph Leiden [133] text cardinality backbone Louvain [92] tweet, URL cardinality,cosine similarity threshold, EDO Louvain [94] tweet, parley, URL, username text similarity, cardinality, cosine similarity threshold, kNN graph Louvain [73] retweet, tweet, URL, hashtag cosine similarity TF-IDF, text similarity threshold, ADO [138, 139] retweet, URL, hashtag, mention, reply cardinality ADJ focal structures [131] retweet, tweet, URL, hashtag, mention cosine similarity TF-IDF ADJ Leiden [102] retweet, hashtag, image, handle change, synchronization Jaccard coefficient, cardinality, cosine similarity threshold, EDO [15] URL cardinality kNN graph Louvain [47, 48, 12, 53, 45, 106] URL cardinality threshold, ADO connected components [2] URL, hashtag, mention unweighted EDO Leiden [91] URL, hashtag, mention cardinality threshold, EDO Louvain [46, 116] URL, text-image cardinality threshold, ADO connected components [96] image cardinality threshold, kNN graph [141] image, video cardinality threshold, ADO connected components [14] hashtag cardinality threshold connected components [89] hashtag cardinality threshold, backbone Louvain [135] hashtag cardinality [58] like cardinality threshold [64] comment cardinality threshold k-means, hierarchical clustering [88] mention cosine similarity TF-IDF threshold Louvain $\dagger$ ADJ: adjacent time window, EDO: evenly distributed overlapping time window, ADO: action-driven overlapping time window
@@
After selecting one or more co-actions to identify similar user behaviors, it is necessary to define the corresponding similarity functions. A similarity function computes the weight $w_{jk}$ of the edges connecting two coordinated users $u_{j}$ and $u_{k}$ based on how similar their behaviors are according to the chosen co-action. Independently of the type of co-action, the majority of works use similarity functions based on the cardinality of the co-action between the two users [14, 15, 18, 51, 59, 63, 64, 68, 94, 96, 93, 95, 92, 109, 110, 139, 138, 32, 135, 47, 48, 12, 53, 45, 46, 116, 141, 137, 121, 19, 27, 133, 106, 52, 66]. For example, when using co-hashtags, the similarity function may simply count the number of times the two users used the same hashtags. Another widely adopted function is the cosine similarity between the two user vectors [28, 73, 94, 92, 99, 20, 56, 70, 29, 102, 122, 131, 28, 88]. These can be binary vectors, frequency vectors, or TF-IDF weighted vectors. The latter allows discounting the importance of popular or viral content, boosting instead the relevance of unpopular items [73, 99, 20, 56, 70, 29, 122, 131, 28, 88]. Depending on the choice of co-action, the previous similarity functions must be preceded by additional processing steps. For example, the use of co-actions such as co-post, co-image, and co-video involve counting the number of times two users shared the same content, which severely limits the possibility to detect certain coordinated behaviors. For this reason, some scholars relaxed this requirement by also considering the posting of similar—as opposed to equal—content. This solution requires defining additional similarity functions for texts, images, and videos. In literature, text similarity was computed via correlation [67], cosine similarity of document embeddings [95, 93, 92, 32, 73, 28], Ratcliff/Obershelp algorithm [101], or Jaccard similarity [52]. All these works first computed similarities between the text of two users’ posts. Then, they set a threshold to select highly similar texts. Finally, they computed user similarities based on the number of similar texts [95, 93, 92, 32, 52], or as the average of the similarities between the highly similar texts [73, 28]. Analogously, co-image involves a pre-processing step for representing the images with their embeddings [96] or with RGB color histograms [102]. Then, image similarity is computed via measures such as the Euclidean distance [96]. Finally, a similarity threshold is applied and user similarity is computed as the cardinality [96] or the Jaccard coefficient [102] of the sets of similar images.
@@
There are three main approaches for coordination network filtering: (i) fixed thresholds, (ii) statistical validation, and (iii) the timings of the co-actions. Methods based on fixed similarity thresholds discard all edges in the network whose weight $w<w_{\text{th}}$, and all the resulting disconnected nodes [14, 18, 58, 63, 64, 73, 95, 93, 96, 89, 91, 92, 101, 102, 110, 24, 134, 32, 47, 48, 12, 53, 45, 46, 116, 141, 137, 19, 66, 88]. The threshold $w_{\text{th}}$ is chosen in such a way to retain only strongly coordinated users, as they are implicitly considered to be more relevant. Relevant nodes can also be identified via eigenvector centrality, by pruning those nodes that do not exceed a certain threshold [73, 28]. The similarity and centrality thresholds are typically selected arbitrarily, without a strong underlying theoretical motivation [99]. Moreover, coordination networks resulting from the analysis of different datasets, or from the use of different types of co-actions, inevitably result in different edge weight and node centrality distributions. This variability makes the repeated use of “standard” thresholds unsuitable and mandates in-depth case-by-case analyses. To alleviate this burden some works leverage $k$-nearest neighbors graphs ($k$-NNG), where two users $u_{j}$ and $u_{k}$ are connected only if $u_{j}$ is among the $k$-nearest neighbors of $u_{k}$ [15, 95, 93, 96]. This filtering operation retains only the $k$ strongest neighborhoods in the coordination network, thus reducing the emphasis on the edge weight. However, determining the best value for $k$ is also challenging since $k$ strongly depends on the characteristics of the different networks and their layers. Other methods for identifying relevant network structures are those that retain statistically-meaningful edges, independently of their weight [44, 112, 127]. This approach, used by several recent works [68, 89, 99, 20, 56, 70, 29, 122, 121, 133], is not biased towards fixed arbitrary levels of similarity or coordination, but instead erases network structures that convey limited information, allowing to focus on meaningful expressions of coordination.
@@
[68, 19] adjacent 1 week user single
@@
[101] action-driven overlapping 10 sec user single
@@
[…]
@@
This step aims to detect a set of coordinated communities $P=\{P_{1},\dots,P_{n}\}$ via community discovery on the coordination network. Multiplex networks are either flattened before performing community discovery [57, 138, 139] or an algorithm suitable for multiplex networks is used [122, 76].
@@
Most of the works dealing with single layer networks carry out community discovery with Louvain [15, 18, 51, 63, 68, 94, 89, 91, 92, 26, 96, 133, 66, 88], or more broadly via modularity clustering [19]. Other recent works [2, 95, 93, 122, 131] relied on Leiden [124]. Among the advantages of these approaches is their scalability, which makes them suitable for the analysis of large networks. In addition, modularity clustering provides a hierarchical community structure, allowing for analyses at different levels of granularity. The authors of [99, 121] combined community discovery via Louvain with the iterative application of a progressively increasing edge weight filtering threshold on the coordination network. As a result of the moving threshold, they studied how the structure and the properties of the coordinated communities change across the whole spectrum of coordination. Moreover, the moving threshold implicitly defines a measure for the extent of coordination observed at each iteration, for each coordinated community, thus providing a continuous score of coordination rather than a binary label. Others proposed a variation of focal structures analysis [111] to identify influential sets of nodes in a coordination network [138, 139]. An alternative to community discovery is the application of a particularly restrictive set of filters that results in splitting the coordination network in multiple connected components [14, 47, 48, 12, 53, 45, 46, 116, 141, 106, 52]. Since each component is disconnected from the others, this process essentially produces an output that is similar to that of community discovery, in that it identifies groups of connected and highly coordinated users. However, the application of very restrictive filters also discards much of the coordination network [99].
@@
[90] URL, hashtag, mention cardinality EDO sum cardinality
@@
The works presented in Table 5 performed community discovery on a multiplex coordination network. Some authors reduced the complexity of dealing with multiplex networks by flattening them into single layer networks where nodes and edges are the union of the nodes and edges of each layer. Flattened networks can be either weighted [90] or unweighted [73, 92, 28], with the latter resulting in the loss of some information. Alternative approaches involve representing the multiplex network as a single-layer multigraph with labeled parallel edges [52], leveraging multi-view clustering to combine different network layers [74, 75], or applying Louvain on each layer and an iterative probabilistic voting consensus algorithm to achieve consensus clustering [32].

### v2

@@ 2.6.1. Taxonomy
Figure 5 demonstrates the generality of the conceptual framework by presenting a taxonomy of coordinated online behavior based on the dimensions of authenticity and harmfulness. This analysis takes inspiration from the well-known conceptual framework of information disorder that categorizes instances of mis-, mal-, and disinformation based on the falseness and harmfulness of the information [143]. As shown in Figure 5, the dimensions of authenticity and harmfulness allow distinguishing between different types of coordinated online behavior, ranging from potentially nefarious to neutral and beneficial ones. Among the most problematic instances of online coordination are harmful and inauthentic phenomena such as information operations and disinformation campaigns, which are often initiated by malicious social bots and state-sponsored trolls [138, 18, 3]. Other problematic phenomena are those featuring harmful yet authentic behaviors, such as the activity of conspiracy theorists and hateful extremists [97, 94, 83]. Conversely, the activity of coordinated crisis communication by democratic institutions during emergencies [151], grassroots social movements, fandoms, and other activists represent instances of harmless and authentic coordination [126]. Finally, harmless yet inauthentic behaviors lay outside of the partially overlapping sets, and are exemplified by hacktivists, dissidents, and other anonymous protesters [126]. Alternative taxonomies can be obtained by leveraging the dimensions of orchestration and time-variance, which would highlight additional phenomena to those shown in Figure 5.
@@ 3. Problem Definition
The problem of identifying and investigating different types of coordinated online behavior involves defining two functions $f(\cdot)$ and $g(\cdot)$ that respectively implement the tasks of coordinated behavior detection and characterization, as outlined in Figure 6. Given a set of users and their actions on one or more online platforms, $f(\cdot)$ identifies possible coordinated groups of users. Instead, $g(\cdot)$ extracts additional information for each detected group, thus contributing to determining the nature, intent, and the overall characteristics of the involved actors (e.g., whether they are inauthentic, harmful, etc.). The detection and characterization tasks are related to the Definition 2.4 of coordinated online behavior and its components in that the function $f(\cdot)$ implementing the detection task does so via the analysis of user actions, while the function $g(\cdot)$ implementing the characterization task provides information about the actors and their intent. third component, i.e., actors and intent, respectively. A comprehensive overview of the entire process is depicted in Figure 6. The input is represented by the set of users $U$ to analyze and their activities $H$. The detection task differentiates coordinated users from non-coordinated ones. Depending on the detection method, the distinction between the two can be expressed as binary labels assigned to the users, as two or more sets (e.g., clusters) of either coordinated or non-coordinated users, or as two or more coordinated or non-coordinated communities (i.e., nodes and edges) from a network. These are subsequently scrutinized during the characterization task, which computes a set of indicators for each coordinated user, set, or community. The indicators are selected so as to provide information about the characteristics of the coordinated actors and their behavior. For example, computing bot scores is a common method to estimate the inauthenticity of coordinated users.
@@ 3.1.2. Methods and output
[…]
@@ 4.1.1. User selection
Multiple choices can be made to select a subset of relevant users. A common choice is to select the most active users as they produce the largest share of actions and content. Most active users can be defined as those who publish a large number of original posts (super-producers[64, 72, 103], or as those having a large number of re-shares (super-spreaders[100, 19, 51, 128, 66, 30, 90, 67]. Other than activity, network centrality or influence, suspicious behavior, location, and timings are used for user selection [70]. Furthermore, many other general criteria can also be adopted, such as selecting all users who posted certain keywords, or all followers of a given user. Combining multiple criteria allows for even more fine-grained user selections.
@@
This step builds a coordination network between the users $U^{\prime}\subseteq U$ previously selected. [^5] A few works follow the general approach of network science methods, but build networks of content rather than users. These are discussed in Section 4.1.5. A coordination network is a type of network where links exist only between coordinated nodes. As per Definition 2.4, coordination implies synergic actions between users. In network science methods, this concept is operationalized with co-actions. A co-action represents two users performing the same action on the same target or content. For instance, two users who comment, like, or re-share the same post are generating a co-action. As shown in Figure 8, the two coordinated users need not be directly connected in the social or interaction network. This characteristic makes coordination networks particularly suitable for surfacing coordination between seemingly unrelated users, such as those involved in inauthentic or harmful behavior, so much so that some scholars specifically refer to latent coordination networks [146]. In the most general case, a coordination network is a multiplex network $G(V,E,W,\mathcal{L})$. In order to build $G$, one must define the types of co-actions $C_{a}$ to consider (e.g., re-shares, mentions, follows, etc.). When multiple co-actions are used, $G$ is a multiplex network with $L$ layers, where $\mathcal{L}=\{1,\dots,L\}$ is the set of layers, each corresponding to a co-action $i\in C_{a}$. However, the majority of existing works leverage a single co-action. In this case, the number of layers is $L=1$ and $G$ is a single layer network. Each layer in $G$ is an undirected weighted graph $G^{i}(V^{i},E^{i},W^{i})$, where $V^{i}\subseteq U^{\prime}$, $E^{i}$ and $W^{i}$ respectively denote the set of nodes, edges, and weights of layer $i$. We highlight that $V=\bigcup_{i\in\mathcal{L}}V^{i}$, $E=\bigcup_{i\in\mathcal{L}}E^{i}$, and $W=\bigcup_{i\in\mathcal{L}}W^{i}$. Given a layer $i$, an edge $e^{i}_{jk}$ is created if there exists a co-action of type $i$ between two users $(u_{j},u_{k})$. The edge weight $w^{i}_{jk}=sim_{i}(u_{j},u_{k})$ is obtained via a similarity function that computes pairwise user similarities in $U^{\prime}$. Different similarity functions can be used for different co-actions.
A: social network $u_{1}$$u_{2}$$u_{3}$
B: interaction network $u_{1}$$u_{2}$$u_{3}$
C: coordination network $u_{1}$$u_{2}$$u_{3}$
@@
reference action similarity filters community detection
[18] retweet cardinality threshold, ADJ modularity clustering
[47] retweet cardinality EDO Louvain
[62, 61] retweet cardinality threshold, ADO Louvain
[65] retweet cardinality backbone, ADJ Louvain
[115] retweet cardinality EDO
[100, 19, 51, 66, 30, 128] retweet cosine similarity TF-IDF backbone Louvain
[129, 114] retweet cosine similarity TF-IDF backbone, EDO Leiden
[144] retweet cardinality threshold, ADJ
[8] retweet cosine similarity kNN graph, correlation HDBSCAN
[17] retweet, tweet cardinality threshold, EDO Louvain, connected components
[59] retweet, tweet cardinality threshold, EDO Louvain
[116] retweet, tweet cardinality threshold, EDO
[13] tweet cosine similarity threshold, EDO
[28] tweet cardinality ADO
[64] tweet cardinality threshold cohesive campaign
[102] tweet text similarity threshold, ADO
[98] parley cardinality threshold, kNN graph Leiden
[140] text cardinality backbone Louvain
[149] tweet threshold, kNN graph, ADO
[95] tweet, URL cardinality,cosine similarity threshold, EDO Louvain
[97] tweet, parley, URL, username text similarity, cardinality, cosine similarity threshold, kNN graph Louvain
[70] retweet, tweet, URL, hashtag cosine similarity TF-IDF, text similarity threshold, ADO
[104] retweet, tweet, URL, hashtag, fast retweet cosine similarity TF-IDF ADO (fast retweet) connected components
[145, 146] retweet, URL, hashtag, mention, reply cardinality ADJ focal structures
[138] retweet, tweet, URL, hashtag, mention cosine similarity TF-IDF ADJ Leiden
[103] retweet, hashtag, image, handle change, synchronization Jaccard coefficient, cardinality, cosine similarity threshold, EDO
[27] retweet, tweet, image, synchronization cosine similarity TF-IDF threshold Louvain
[67] interaction, text, synchronization Normalized Compression Distance kNN graph
[71] text, synchronization, hashtag, URL, duet, stitch, reply cosine similarity TF-IDF threshold connected components
[148] hashtag, URL, video description, music, audio cardinality kNN graph connected components
[20] URL cosine similarity TF-IDF threshold, EDO connected components
[16] URL cardinality kNN graph Louvain
[43, 44, 14, 49, 41, 110, 112, 122] URL cardinality threshold, ADO connected components
[150] URL cardinality
[3] URL, hashtag, mention unweighted EDO Leiden
[94] URL, hashtag, mention cardinality threshold, EDO Louvain
[22] URL, text cosine similarity TF-IDF threshold, ADJ
[42, 121] URL, text-image cardinality threshold, ADO connected components
[99] image cardinality threshold, kNN graph
[152] image, video cardinality threshold, ADO connected components
[15] hashtag cardinality threshold connected components
[91] hashtag cardinality threshold, backbone Louvain
[142] hashtag cardinality
[54] like cardinality threshold
[60] comment cardinality threshold k-means, hierarchical clustering
[106] comment cardinality threshold, ADO connected components
[90] mention cosine similarity TF-IDF threshold Louvain
[157] report cardinality threshold connected components
[158] follow cosine similarity Louvain
$\dagger$ ADJ: adjacent time window, EDO: evenly distributed overlapping time window, ADO: action-driven overlapping time window
@@
There are three main approaches for coordination network filtering: (i) fixed thresholds, (ii) statistical validation, and (iii) the timings of the co-actions. Methods based on fixed similarity thresholds discard all edges in the network whose weight $w<w_{\text{th}}$, and all the resulting disconnected nodes [15, 17, 54, 59, 60, 70, 98, 99, 91, 94, 95, 102, 103, 116, 24, 141, 32, 43, 44, 14, 49, 41, 42, 121, 152, 144, 18, 62, 90, 13, 149, 71, 20, 22, 106, 157, 92, 93, 81, 61]. The threshold $w_{\text{th}}$ is chosen in such a way to retain only strongly coordinated users, as they are implicitly considered to be more relevant. Relevant nodes can also be identified via eigenvector centrality, by pruning those nodes that do not exceed a certain threshold [70, 29]. The similarity and centrality thresholds are typically selected arbitrarily, without a strong underlying theoretical motivation [100]. Moreover, coordination networks resulting from the analysis of different datasets, or from the use of different types of co-actions, inevitably result in different edge weight and node centrality distributions. This variability makes the repeated use of “standard” thresholds unsuitable and mandates in-depth case-by-case analyses.
To alleviate this burden some works leverage $k$-nearest neighbors graphs ($k$-NNG), where two users $u_{j}$ and $u_{k}$ are connected only if $u_{j}$ is among the $k$-nearest neighbors of $u_{k}$ [16, 98, 99, 8, 149, 67, 148]. This filtering operation retains only the $k$ strongest neighborhoods in the coordination network, thus reducing the emphasis on the edge weight. However, determining the best value for $k$ is also challenging since $k$ strongly depends on the characteristics of the different networks and their layers. Other methods for identifying relevant network structures are those that retain statistically-meaningful edges, independently of their weight. This approach, used by several recent works [65, 91, 100, 19, 51, 66, 30, 129, 128, 140], is not biased towards fixed arbitrary levels of similarity or coordination, but instead erases network structures that convey limited information, allowing to focus on meaningful expressions of coordination.
@@
[65, 18, 22] adjacent 1 week user single
[1] adjacent 1 hour, 1 day, content single
@@
[20, 71] evenly distributed overlapping 1 day user single
[13] evenly distributed overlapping 2 day user single
[81] evenly distributed overlapping 6 hour, 1 week user multiple (L=5)
@@
[102, 104] action-driven overlapping 10 sec user single
@@
[149] action-driven overlapping 1 hour user single
@@
Filtering methods based on the timings of the co-actions use a sequence of time windows of equal size $Z=t_{\text{end}}-t_{\text{start}}$. The filtering typically occurs when computing user similarities, by only retaining the co-actions that occur within the same time window. In other words, this filter corresponds to adding a further constraint to the actions that determine the coordination, in that such actions must be temporally close to one another. Table 3 displays the type and length of the time windows used in literature, along with the type of coordination network built. As shown in Tables 3 and 4, time windows can be either adjacent [3, 65, 138, 145, 146, 18, 1] or overlapping. Furthermore, overlapping time windows can be evenly distributed in time [47, 17, 59, 115, 116, 94, 3, 72, 95, 48, 73, 93, 103, 129, 20, 71, 13, 81], or action-driven—that is, positioned based on the timings of the actions [62, 28, 102, 29, 70, 41, 49, 41, 43, 44, 42, 121, 110, 14, 152, 32, 141, 104, 149, 61]. As shown in Table 4, in the latter case each time window starts when a relevant action occurs. In the former case instead, the additional parameter step $\delta<Z$ defines the temporal offset between two consecutive time windows. The amount of overlap $O$ between two consecutive overlapping time windows is thus $O=Z-\delta$. Given a sequence of actions by two or more users, the choice of time windows and their parameters influence the number of such actions that are valid co-actions. These are indicated with the lock icon in the example shown in the rightmost column of Table 4. Let $d=t_{2}-t_{1}$ be the delay with which user $u_{2}$ performs an action at time $t_{2}$, with respect to the same action that $u_{1}$ performed at time $t_{1}$. Independently of the type of time window, actions whose $d>Z$ are never considered as co-actions. Then, actions with $d\leq Z$ are always co-actions when using action-driven overlapping time windows. Conversely, actions with $d\leq O$ are always co-actions when using evenly distributed overlapping time windows, and can possibly be co-actions when $O<d\leq Z$. Instead, with adjacent time windows, also actions that are close in time (i.e., $d\ll Z$) are occasionally not considered valid co-actions, when they occur across the boundary between two windows, as in the case of the actions occurred at $t_{3}$ and $t_{4}$ in the example of Table 4. Therefore, using overlapping rather than adjacent time windows ensures that no actions close in time are missed. However, it increases the number of windows required to cover the same time frame, leading to more computational demands. Another critical consideration is the length $Z$ of the time window. A larger $Z$ includes more actions as co-actions, while a shorter one restricts valid co-actions to highly synchronized events, while also increasing the number of windows and the resulting computations. The choice of $Z$ also depends on the goal of the analysis. Literature indicates that inauthentic or harmful coordinated behaviors are highly synchronized and can often be detected with short windows [103, 73], whereas emergent human behaviors that are typically less orchestrated and loosely synchronized, require longer windows to capture their relaxed temporal dynamics [100].
Table 4. Time window types, their parameters, and their effect on valid co-actions. To the right, a sequence of actions occurs at times $t_{1},\ldots,t_{6}$. The same sequence results in different valid co-actions, marked by the lock icon, depending on the time window type.
@@
This step aims to detect a set of coordinated communities $P=\{P_{1},\dots,P_{n}\}$ via community discovery on the coordination network. Multiplex networks are either flattened before performing community discovery [145, 146] or an algorithm suitable for multiplex networks is used [129, 74].
@@
Most of the works dealing with single layer networks carry out community discovery with Louvain [16, 17, 47, 59, 65, 97, 91, 94, 95, 26, 99, 140, 62, 90, 158, 61], or more broadly via modularity clustering [18]. Other recent works relied on Leiden [3, 98, 129, 138, 114]. Among the advantages of these approaches is their scalability, which makes them suitable for the analysis of large networks. In addition, modularity clustering provides a hierarchical community structure, allowing for analyses at different levels of granularity. The authors of [100, 128] combined community discovery via Louvain with the iterative application of a progressively increasing edge weight filtering threshold on the coordination network. As a result of the moving threshold, they studied how the structure and the properties of the coordinated communities change across the whole spectrum of coordination. Moreover, the moving threshold implicitly defines a measure for the extent of coordination observed at each iteration, for each coordinated community, thus providing a continuous score of coordination rather than a binary label. Others proposed a variation of focal structures analysis to identify influential sets of nodes in a coordination network [145, 146].
An alternative to community discovery is the application of a particularly restrictive set of filters that results in splitting the coordination network in multiple connected components [15, 43, 44, 14, 49, 41, 42, 121, 152, 110, 48, 104, 148, 20, 112, 122, 157, 106]. Since each component is disconnected from the others, this process essentially produces an output that is similar to that of community discovery, in that it identifies groups of connected and highly coordinated users. However, the application of very restrictive filters also discards much of the coordination network [100].
@@
[52] retweet, URL, hashtag, mention temporal weighted cardinality Generalized Leiden
@@
[93, 92] URL, hashtag, mention cardinality EDO sum cardinality
[81] URL, hashtag, mention, reply, retweet cosine similarity TF-IDF threshold, EDO Generalized Louvain, Generalized Infomap
@@
The works presented in Table 5 performed community discovery on a multiplex coordination network. Some authors reduced the complexity of dealing with multiplex networks by flattening them into single layer networks where nodes and edges are the union of the nodes and edges of each layer. Flattened networks can be either weighted [93] or unweighted [70, 95, 29], with the latter resulting in the loss of some information. Alternative approaches involve representing the multiplex network as a single-layer multigraph with labeled parallel edges [48], leveraging multi-view clustering to combine different network layers [72, 73], or applying Louvain on each layer and an iterative probabilistic voting consensus algorithm to achieve consensus clustering [32]. Only a couple of works exploit the multiplex version of Leiden [52], Infomap [81], and Louvain [81], fully exploiting the multimodal structure of the coordination network.